Three kilograms of neon in a constant volume system. The neon is initially at 550 kPa and 350 K. Its pressure is increased to 2000 kPa by a paddle work with 210 kJ of heat added. Find (a) the final temperature, (b) the change in internal energy, (c) the change in enthalpy, (d) the nonflow work. 


1
Expert's answer
2021-12-06T09:40:20-0500

"m = 3 \;kg \\

p_1 = 550 \; kPa \\

p_2 = 2000 \; kPa \\

T_1 = 350 \;K \\

Q_{in} = 210 \;kJ"

(a) The final temperature

"Volume = constant \\

pV = mRT \\

\frac{p}{T} = \frac{mR}{T} = constant \\

\frac{p_1}{T_1} = \frac{p_2}{T_2} \\

T_2 = \frac{p_2T_1}{p_1} \\

T_2 = \frac{2000 \times 350}{550} = 1272.73 \;K"

(b) From table Ideal gas specific heat for neon:

"c_p = 1.0299 \; kJ/kg \;K \\

R = 0.4119 \;kJ/kg \;K \\

c_p-c_v =R \\

1.0299 -c_v = 0.4119 \\

c_v = 0.618 \;kJ/kg \;K"

Change in internal energy

"ΔU =mc_v(T_2-T_1) \\

ΔU = 3 \times 0.618(1272.73-350) = 1710.74 \; kJ"

(c) Change in enthalpy

dH = ΔU + pdV

dH = ΔU = 1710.74 kJ

(d) The nonflow work "= \int pdV = \int p(V_2-V_1) = 0"

"W_{nonflow} = 0"